OpenStudy (nikykot):

Using Euler’s Formula (F + V = E + 2), find the number of vertices. Edges: 33 Faces: 18 Vertices: ? A. 17 B. 18 C. 20 D. 16

OpenStudy (owen3):

This is an algebra problem. Did you plug in values?

OpenStudy (owen3):

(F + V = E + 2), find the number of vertices. Edges: 33 Faces: 18 Plugging in 18 + V = 33 + 2 can you solve it now?

OpenStudy (nikykot):

I got 53...how do you I find the vertices?

OpenStudy (nikykot):

@owen3

OpenStudy (nikykot):

oh never mind I get it know, thank you. The answer I got was 17